A solid weighs 0.040N in air and 0.024N when fully immersed in a liquid of density
800kgm. Calculate the
volume of the solid (g = 10ms2)
(a) 2.0 x 10-6m3 (b) 2.5 x 10-6m3 (c) 3.0 x10-6m3 (d) 2.0 x 10-6m3
1
Expert's answer
2019-06-21T09:15:08-0400
"W_a-W_l=\rho g V"

The volume of the solid:

"V=\frac{W_a-W_l}{\rho g}"

"V=\frac{0.04-0.024}{(800)(10)}= 2.0 \cdot 10^{-6}m^3"